OpenStudy (anonymous):

portable DVD player requires six 1.5-volt batteries to operate. If the resistance in the circuit is 7.2 ohms, what is the current? Select one of the options below as your answer: A. 0.21 amps B. 1.3 amps C. 11 amps D. 65 amps E. 1.8 amps

OpenStudy (fifciol):

I=V/R=6*1,6/7.2=1,25 A, round it off to 1,3 Answer B

OpenStudy (anonymous):

SWAG^

OpenStudy (anonymous):

yes^^